OpenStudy (anonymous):

solve for x by completing the square 1. x^2 - mx - 6m^2 = 0 2. kx^2 + px + q = 0

OpenStudy (amoodarya):

\[x^2-mx-6m^2=\\(x-\frac{ m }{ 2 })^2-(\frac{ m }{ 2 })^2-6m^2=0\]